OpenStudy (anonymous):

will give medal 3 questions help please, 1. Suppose y varies inversely with x. Write an equation for the inverse variation. Y = 5 when x = -5 2. Graph each inverse variation. Y = -18 / x 3.Identify the excluded value for each function f(x)=5/x

OpenStudy (anonymous):

1. y = k/x 4 = k/-6 k = -24 y = -24/x x = -24/y

OpenStudy (anonymous):

are you sure this is current ?

OpenStudy (anonymous):

most likely yes

OpenStudy (anonymous):

can you do the others please

OpenStudy (anonymous):

3. Excluded values occur to prohibit division by 0 (zero). f(x) = 5/x then this function is meaningless when x = 0 and it is said to be discontinuous at x = 0. Take your denominator and set it not equal to zero, since dividing by zero would give you an answer that is undefined. In the case of this question the denominator is x, therefore x=/=0, since a x value equal to 0 would cause the denominator to be zero.

OpenStudy (anonymous):

okay thank you so much what about number 2 ?

OpenStudy (anonymous):

sorry can't help u with that

OpenStudy (tester97):

For number 2 just plug that into your graphing calc if you dont have one use wolframalpha :)

OpenStudy (anonymous):

@tester97 your right i totally forgot about wolfram lolll thanks

OpenStudy (tester97):

I did it for you already so here ya go enjoy :) http://www.wolframalpha.com/input/?i=Y+%3D+-18+%2F+x

OpenStudy (anonymous):

@ryanvarghese12790 i have two more Identify the asymptotes of the graph of each function. Then graph the function. y=1/x+2 Identify the asymptotes of the graph of each function. Then graph the function. Y = 5 / x – 4 + 1

OpenStudy (anonymous):

omgggg @tester97 thank you thats correct i wish i could give you a medal

OpenStudy (tester97):

welcome |:) just use wolframalpha to graph these last two ^_^

OpenStudy (anonymous):

ohhhh your right i can well then i have a word problem if you guys can help

OpenStudy (anonymous):

sorry can't help u with the other two q's, this is not one of my strongest topics, I am terribly sorry for the inconvenience this has caused @bettylee

OpenStudy (anonymous):

its okay you helped me with the others which i very much appreciated loll you dont have to apologized you actually helped a lot. but here is a word problem 1.for a 225-watt bulb, the intensity i of light of humens at a distance of x feet is i=225/x^2 A.what is the intensity of light 5ft from the bulb B. suppose your distance from the bulb doubles, how does the intensity of the light change ? explain

OpenStudy (tester97):

>.< sorry im not good with word probz

OpenStudy (anonymous):

okay loll well i wasd talking to ryan lolll hopefully he can help, ummm can you graph this on wolfram for me Y = 5 / x – 4 + 1

OpenStudy (anonymous):

thankssss @tester97 you rock! ( :